OpenStudy (anonymous):

\[(x ^{-5} \div y ^{-8})^{4}\] I want to have a detailed explanation of the solution. Been struggling to understand the rules that apply on each step. I am pretty thorough with the product rules and everything else but I have hard time with the ones where the base is different. Any sort of help is appreciated here!

OpenStudy (kaiz122):

the equation can also be written as \[\frac{(x^{-5})^4}{(y ^{-8})^4}\]

OpenStudy (kaiz122):

we have \[(n^a)^b = n^{a \times b}\] try it both for the numerator and denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what I get confused with is that how should I reciprocate? One of the rules I read was to reciprocate the values when there is a negative sign. So two questions really... 1) Do I need to reciprocate? If yes, then Will the reciprocation invert the sign outside the bracket or just the sign of the denominator? 2) How is the reciprocation done step-by-step?

OpenStudy (kaiz122):

yes that was true, if any values that have a negative exponent will be reciprocate, in this case, you can either reciprocate it first, or multiply it first by the exponents and then reciprocate.

OpenStudy (anonymous):

\[x ^{-20}\div y ^{-32}\]

OpenStudy (kaiz122):

yes, and then get the reciprocal of the numerator and denominator

OpenStudy (kaiz122):

what would be the answer @sunnymsus ?

OpenStudy (anonymous):

k...would it be \[y ^{32} \div x ^{20}\]

OpenStudy (kaiz122):

yes it is. :)

OpenStudy (anonymous):

o nice...that was easy. let me try another problem like that.

OpenStudy (anonymous):

another thing...what if the numerator was positive and the denominator was negative?

OpenStudy (kaiz122):

if you always practice, it would be a lot easier for you.

OpenStudy (anonymous):

yes...i have been trying since yesterday...although i have become better at it but it's still not perfect. hoping with practice i would be able to crack it.

OpenStudy (kaiz122):

Good job! :)

OpenStudy (anonymous):

what if the numerator was positive and the denominator was negative?

OpenStudy (anonymous):

will we keep the sign of the numerator as positive even after reciprocation?

OpenStudy (anonymous):

the site where I am practicing says the signs are off. It says it is incorrect.

OpenStudy (kaiz122):

if the numerator has positive exponent and the denominator has negative, only reciprocate the denominator @sunnymsus what's the problem?

OpenStudy (anonymous):

if you pay attention to the answer I had pasted above, I changed the sign's of both numerator and denominator, whereas I think I should have done that only for the denominator? Is that correct?

OpenStudy (kaiz122):

the answer above is correct. since both have negative exponents we will take their reciprocals.

OpenStudy (kaiz122):